OpenStudy (anonymous):

An architect for a golf course wants to plan a sand trap that passes between a tree and a cart path. Using these as the focus and directrix, how can the architect plan a parabolic sand trap that will be equidistant from the tree and the cart path at all times? Describe your method in full sentences.

OpenStudy (anonymous):

@morganmay19

OpenStudy (morganmay19):

idk

OpenStudy (anonymous):

xc ok...

OpenStudy (anonymous):

@HelpBlahBlahBlah

OpenStudy (anonymous):

I'm sooooo bad at parbola's and this stuff, hang on.

OpenStudy (anonymous):

I'm sorry I have no idea :/

OpenStudy (anonymous):

OHHMAIIIGERDD!

OpenStudy (anonymous):

have you tried google?

OpenStudy (anonymous):

No, I want to know how to do it, not just the answer.

OpenStudy (anonymous):

okay hold on

OpenStudy (anonymous):

The trap should coincide with the curve of the parabola, the vertex of which would be half between the tree and the cart path.The equation would be given as y= a(x-h)^2+k, where (h,k) is the vertex and (x,y)is the coordinate of any point on the parabola. If the focus is at (Fx,Fy) and the direction is y, p=1/2(Fx-y) and a=1 / 4p or a=1/4[1/2(Fx-y)] a=1/2(Fx-y) so The equation of the sand trap is given by y=[1/2(Fx-y)] (x-h)^2+k

OpenStudy (anonymous):

Thanks a lot @lexikaylynn xD

OpenStudy (anonymous):

your welcome :D